what is 3/5(2/3x-5/6)=3
Answer by HyperBrain(694)

3/5(2/3x-5/6)=3
Multiply 5 to both sides
3(2x/3-5/6)=5*3
Divide both sides by 3
2x/3-5/6=5
Multiply both sides by 6
4x-5=30
4x=35
x=35/4
Power up,
HyperBrain!
Stop this mess! http://uselessplastic.wordpress.com/
http://uselessplastic.multiply.com/